Consider the surface S = n (x, y, z) | z = p x 2 + y 2 and 1 ≤ z ≤ 3 o .(a) Sketch the surface S in R 3 . Also show its XY-projection on your sketch. (2) (b) Evaluate the area of S, using a surface integral


1
Expert's answer
2022-01-24T15:54:47-0500

a)

"z=\sqrt{x^2+y^2}"




XY-projection:




The XY-projection the sketch is the shaded region


b)

"x=zcos\varphi"

"y=zsin\varphi"

"z=z"

"r(z,\varphi)=zcos\varphi i+zsin\varphi j+zk"

"r_z=cos\varphi i+sin\varphi j+k"

"r_{\varphi}=-zsin\varphi i+zcos\varphi j"

"r_z\times r_{\varphi}=\begin{vmatrix}
 i & j&k \\
 cos\varphi & sin\varphi &1\\
-zsin\varphi &zcos\varphi &0
\end{vmatrix}=-zcos\varphi i -zsin\varphi j+zk"


"|r_z\times r_{\varphi}|=z\sqrt 2"


surface area:

"S=\iint |r_z\times r_{\varphi}|dA=\sqrt 2\int ^{2\pi}_0\int^3_1 zdzd\varphi =8/\sqrt 2\int ^{2\pi}_0 d\varphi=8\pi\sqrt 2"